### v1

@@ Abstract.
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\tfrac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even. This extends the validity of the Fan–Yu–Wang bound from the range $3\leq\chi\leq\frac{n}{2}$ to the full range $3\leq\chi\leq n-1$.
@@ 1. Introduction
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\frac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\frac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
@@ 3. Main results
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\frac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\frac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
@@
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\frac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\frac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
@@ 4. A comparison with Wilf’s bound
Indeed, take $G=K_{3,3}$. Here $n=6$ and the least adjacency eigenvalue is $\lambda_{n}(G)=-3$. Plugging $\lambda=-3$ and $n=6$ into the right-hand side of (3.12) gives
@@ Acknowledgements
The authors would like to thank Prof. Jie Ma for helpful comments and suggestions.

### v2

@@ Abstract.
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)\vee\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even. This extends the validity of the Fan–Yu–Wang bound from the range $3\leq\chi\leq\frac{n}{2}$ to the full range $3\leq\chi\leq n-1$.
@@ 1. Introduction
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)\vee\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
@@ 3. Main results
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)\vee\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
@@
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)\vee\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
@@ 4. A comparison with Wilf’s bound
When $\chi=2$, Theorem 3.1 simplifies to $\lambda_{n}\geq-n/2$, which was proved by Constantine [1] for all graphs. Theorem 3.2 is therefore true when $\chi=2$, and we can let $G=K_{3,3}$ for which $n=6$ and the least adjacency eigenvalue is $\lambda_{n}(G)=-3$. Plugging $\lambda=-3$ and $n=6$ into the right-hand side of (3.12) gives
@@ Acknowledgements
The authors would like to thank Prof. Jie Ma for helpful comments and suggestions. The authors also thank Hitesh Kumar for pointing out several typos in an earlier version of this paper.